ratio between boys to girls is 7:5 and there are total 60 students.how many more boys are there than girls?
Answer by tommyt3rd(5050)

b/g=7/5 so that b=7/5 g
and
b+g=60 so that 7/5 g + g = 60

g=25 (so b=35)
that means there are 10 more boys than girls